# Model-independent unbinned analysis of $B\to K^{*}(\to K^{+}\pi^{-})\mu^{+}\mu^{-}$: zeroes, bounds, Wilson coefficients and symmetries

### v1

@@ I Introduction
The purpose of these analyses was to extract Wilson coefficients together with hadronic information directly from the data using an Effective Field Theory (EFT)-based functional form for the transversity amplitudes. These analyses differ in the way that the hadronic contributions are parameterised. On the one hand, Ref. Aaij et al. (2024a) relies on the use of a $z$-expansion to parameterise the non-local charm-loop contributions and includes constraints from $B(B\to K^{*}J/\psi)$ and $B(B\to K^{*}\psi(2s))$ decays. In this analysis, two hypotheses are considered for the data included in the fit: (i) $q^{2}<0$, where constraints coming from the light-cone Operator Product Expansion (OPE) calculation in Ref. Gubernari et al. (2021) are included; and (ii) $q^{2}>0$, where these constraints are excluded. By contrast, the alternative analysis of Ref. Aaij et al. (2024b) uses a dispersive approach based on the resonance model of Ref. Cornella et al. (2020) and data from the decay $B\to K^{*}\mu^{+}\mu^{-}$ alone. An interesting outcome of both of these analyses is that the data indicate that the charm-loop is insufficient to explain the observed anomalies, and a new physics contribution is required.
@@ II SM predictions for the unbinned observables
While the contributions from local form factors are included in $\mathcal{A}_{V,A}^{\mu}$, non-local charm-quark loop contributions appear in the term $\mathcal{H}^{\mu}$. These effects are accounted for by recasting the phenomenological parameterisation provided in Eq. (7.14) of Ref. Khodjamirian et al. (2010). This parameterisation is itself based on the light-cone OPE and LCSR calculations of the non-local form factor in the region $q^{2}\ll 4m_{c}^{2}$, and its corresponding matching to the dispersion relation proposed in the same paper. In practice, this is implemented as a shift in the factorisable charm-quark loop function $Y(q^{2})$ Buras et al. (1994)
@@
where $s_{0}=1$ GeV2. The parameters $a^{\perp,\parallel,0}$, $b^{\perp,\parallel,0}$, and $c^{\perp,\parallel,0}$ are determined by ensuring that they reproduce the phenomenological model of Ref. Khodjamirian et al. (2010) in the $q^{2}$ range from 1 $1\text{ GeV}^{2}$ to 9 $9\text{ GeV}^{2}$. Specific values for these coefficients can be found in Ref. Descotes-Genon et al. (2016).
@@ II.1 Definition of unbinned observables
The uncertainties for the SM predictions are computed using a Monte Carlo sampling approach, which assumes a multivariate Gaussian distribution for the nuisance parameters described earlier, along with variations in the renormalisation scale, masses, and other relevant factors. The angular observables gives rise to non-linear behaviour, particularly stemming from the normalisation terms. A Gaussian scan of the parameters can therefore lead to non-Gaussian distributions of the observables. This is particularly the case when the normalisation is close to zero. To extract meaningful uncertainties we apply a rescaling factor of $1/3$ to linearise the uncertainty propagation.
@@
The SM prediction for the optimised observables $P_{i}$ and $F_{L}$ are shown graphically in Figs. 1 and 2. The $P_{i}^{\rm CP}$ are, to a high accuracy, very close to zero and we do not include them here. For completeness we also provide the non-optimised observables $S_{i}$ and the branching fraction in Appendix A. The unbinned definitions of the $S_{i}$ observables can be easily derived following the same procedure as in Eq. 12. Some remarks are in order here:
@@
The next issue to address is how to approach unbinned data. Here there are two main options: i) either one parameterises the amplitudes and computes the observables from these amplitudes or ii) one determines the observables directly. In the former case, the observables depend on the model used to parameterise the amplitudes. For example, whether the contribution from NP scalars, tensors or additional operators beyond the SM are included, the modelling of hadronic contributions and the treatment of $m_{K\pi}$ and $q^{2}$. There are different analyses in the literature parameterising the amplitudes Aaij et al. (2024a, b). Instead, the latter case, determining observables, is model-independent. In this way we extract the information without hypothesis a priori and we can perform several tests on the self-consistency of the analysis. In the following subsections II.2 and II.3 we define the tests that a consistent theoretical model of the amplitudes should fulfill. In section III, we describe a model-independent method to determine unbinned observables using a moments approach.
@@ II.2 Testing a theoretical model of the amplitudes: set of conditions
If a parameterisation of the amplitudes does not respect the conditions, for instance, by fixing some degrees of freedom with theory in a way that is inconsistent with the data, this can manifest itself as the inability to reproduce one or more observables. In order to remedy this, more degrees of freedom must be included or some of the fixed degrees of freedom must be allowed to vary.
@@
In the presence of scalar NP, the last two conditions in Eq. LABEL:nJ are affected and also the last condition in Eq. 17. In both cases, the left hand side (LHS) of these conditions has to be modified. In a similar way, one can write a corresponding set of equations for the CP conjugate $J$ and $n_{i}$.
@@ II.3 Interpretation of conditions
In order to test a given amplitude parameterisation using the conditions in Eq. II.2 and Eq. 17, on the LHS of these expressions one should insert the parameterisation of the amplitudes as a function of the relevant free parameters (phases and moduli); and on the RHS, experimental values for the coefficients $J$. A fit for the free parameters can then be used to check if it is possible to correctly reproduce the RHS of each equation.
@@
    A fit to data using only $q^{2}>0$ information but adding lattice calculations. The fit for the observables in Ref. Aaij et al. (2024a) reproduces the binned data very well. We therefore take this fit for the observables as the “data”, i.e., the RHS of the conditions in Eq. II.2 and Eq. 17. The results of this fit are depicted in blue in Fig. 3 and in Fig. 15 in Appendix. B.
@@
    A fit to data using, in addition, the $q^{2}<0$ information. This fit is mainly dominated by theory and we reinterpret it as the LHS of the conditions. The results of this fit are depicted in red in Fig. 3 and in Fig. 15.
@@
The problem observed in condition 4 is linked to the problem observed in the comparison between ${\rm Im}A_{\|}$ for $q^{2}>0$ and $q^{2}<0$ in Ref. Aaij et al. (2024a). To a much lesser extent there is also a problem in ${\rm Im}A_{\perp}$, as discussed below.
@@
${\cal A}_{\lambda}^{L,R}$ $=$ ${\cal N}_{\lambda}\left(\left[C_{9}\mp C_{10}\right]{\cal F}_{\lambda}(q^{2})+\frac{2m_{b}M_{B}}{q^{2}}\left[{\cal F}_{\lambda}^{T}(q^{2})\right.\right.$ (22)
$\left.\left.-16\pi^{2}\frac{M_{B}}{m_{b}}{\cal H}_{\lambda}]\right)\right)$
@@ III A new method to determine unbinned observables: A “continuous” moments method
Here we introduce a novel method to extract the full set of model-independent angular observables continuously in $q^{2}$ and $m_{K\pi}$. We do this by extending the previous binned method-of-moments technique into an unbinned one. This is achieved with the use of a $k$-nearest neighbours (knn) algorithm Cover and Hart (1967). The knn with uniform weighting is exactly the definition of the normalised moment (the moment divided by the total weight), $M_{i}$, but instead of being binned, it is really a reflection of the density of the data in the locale of a given phase-space point. Thus one can pick an arbitrary point in $q^{2}$, $m_{K\pi}$, or $(q^{2},m_{K\pi})$, and estimate the value of the angular moment there.
@@
With this unbinned method, different points in the phase space are statistically correlated with each other. The degree of correlation depends on the number of neighbours in the knn, the smoothing and how far apart the points are. When using the results of this method in further analysis for the underlying physics parameters, knowing these correlations is vital. Consequently, if the correlations are well understood, the precise choice of the number of neighbours or smoothing strength is to some extent a matter of aesthetics. Few neighbours leads to a large apparent uncertainty at a single point but nearby points are relatively independent and so give equal statistical power. Conversely, a large number of neighbours reduces the apparent uncertainty but leads to large statistical correlations with neighbouring regions. What really matters is choosing the analysis parameters that maximise the sensitivity to what is being measured, and in particular allows for the resolution of the phase-space dependence of the observables. Care must be taken, as using too many neighbours can lead to a poor description of the true variables in regions where they vary quickly. However, for some analyses, such as finding the zero-crossing points in $q^{2}$ (discussed in the following sections), one may choose a larger number of neighbours, as the zeros lie in a $q^{2}$ region where the observables of interest are expected to vary smoothly.
@@
Previous experimental analyses have assumed that the leptons are massless, reducing the number of angular terms to be measured and so simplifying the task. However, it is found that the effects of lepton mass in the muon mode are non-negligible even at relatively large values of $q^{2}$ ($\sim 4\,\,\text{GeV}^{2}$), as shown in Fig. 5. In the absence of S-wave contributions, the full set of angular observables may be extracted with the unbinned moments, fully accounting for the effects of the lepton masses. The observables $S_{1}^{s}$, $S_{1}^{c}$, $S_{2}^{s}$, $S_{2}^{c}$ do not have orthogonal angular functions and must be extracted from a linear combination of the corresponding moments
@@ IV.1.1 Geometrical tests associated with the structure of the observables
It is important to notice when comparing to data that the LHCb collaboration use a different convention for the observables than the original definition in Ref. Descotes-Genon et al. (2013b). These differing conventions can be interchanged using,
@@
The terms $\omega_{r,i}$ contain the scalar contributions. If the signs of $P_{5}^{\prime}$ and $\omega_{r}$ are the same the bound in Eq. 39 may be violated in an analysis measuring observables, indicating the presence of a scalar NP contribution according to Eq. 47. Similarly, if the signs of $P_{6}^{\prime}$ and $\omega_{i}$ are the same, the bound in Eq. 35 may also be violated according to Eq. 48. All other bounds are unaffected by the presence of a scalar.
@@ IV.1.2 Tests based on the non-trivial relation between the observables
At this point there are two options: 1) to test directly the relation in Eq. 52, which is given in terms of $J_{i}$; or 2) to re-express it in terms of the $P_{i}$ and $P_{i}^{CP}$ observables, as shown below. The latter is implemented by inverting the definitions in Eq. 12.
There is a minor difference between the two options, because in the first case a breaking of Eq. 52 can only be due to an internal problem of the experimental analysis, or to a large new scalar contribution. In the second case, if we use the expression in terms of $P_{i}$ and $P_{i}^{CP}$ and then assume that all $P_{i}^{CP}$ are negligible, we introduce a third possible source of breaking if there exists large NP weak phases in Nature. In any case, this relation is always a very efficient cross-check of an experimental analysis.
We have explicitly checked that, up to machine precision, the relation of Eq. 52 using the $J$’s obtained from the fit of Ref. Aaij et al. (2024a) is fulfilled, considering either the $q^{2}<0$ or $q^{2}>0$ cases. Since the $J_{i}$ are obtained from an amplitude analysis where no scalar contribution is introduced, the relation in Eq. 52 must, by construction, be fulfilled. However, this implies nothing about the presence or absence of a scalar amplitude.
In order to express Eq. 52 in terms of optimised observables we use the definitions of these observables in Eq. 12. One can then write down one relation for $\bar{P_{i}}=P_{i}+P_{i}^{\rm CP}$ and one for $\hat{P_{i}}=P_{i}-P_{i}^{\rm CP}$. Taking the example of $P_{2}$, we can write an equation for $\bar{P}_{2}$ and one for $\hat{P}_{2}$ as a function of all the other observables. The corresponding equations are Matias and Serra (2014)
${P_{2}+P_{2}^{\rm CP}}=+\frac{1}{2{\bar{k}_{1}}}\bigg{[}({\bar{P}_{4}^{\prime}}{\bar{P}_{5}^{\prime}}+\delta_{1})+\frac{1}{\beta}\sqrt{(-1+{\bar{P}_{1}}+{\bar{P}_{4}^{\prime 2}})(-1-{\bar{P}_{1}}+\beta^{2}{\bar{P}_{5}^{\prime 2}})+\delta_{2}+\delta_{3}{\bar{P}_{1}}+\delta_{4}{\bar{P}_{1}}^{2}}\bigg{]},$ (53)
@@
${P_{2}-P_{2}^{\rm CP}}=+\frac{1}{2{\hat{k}_{1}}}\bigg{[}({\hat{P}_{4}^{\prime}}{\hat{P}_{5}^{\prime}}+\hat{\delta}_{1})+\frac{1}{\beta}\sqrt{(-1+{\hat{P}_{1}}+{\hat{P}_{4}^{\prime 2}})(-1-{\hat{P}_{1}}+\beta^{2}{\hat{P}_{5}^{\prime 2}})+\hat{\delta}_{2}+\hat{\delta}_{3}{\bar{P}_{1}}+\hat{\delta}_{4}{\hat{P}_{1}}^{2}}\bigg{]},$ (54)
@@
where $\Omega=\frac{1}{\beta}\sqrt{(-1+P_{1}+{P_{4}^{\prime}}^{2})(-1-P_{1}+\beta^{2}{P_{5}^{\prime}}^{2}})$. For illustration, we show in Fig. 11 the expression in Eq. 55 rewritten in quadratic form using the results from one pseudo-experiment.
As discussed at the beginning of this subsection IV.1, in the presence of a NP scalar the non-trivial relation between the observables is broken. Here we estimate, for the first time, the amount of breaking of the relation due to the presence of a scalar by combining theory inputs with data on the branching fraction of ${B_{s}\to\mu^{+}\mu^{-}}$. If we allow for the presence of a scalar, Eq. 53 and Eq. 54 can be generalised by introducing the following substitutions in Eq. 53
@@
are the CP counterparts of the $\omega_{r,i}$ defined in Eq. 49 and Eq. 50, with $N$ given by Eq. 51. By expanding the terms in $\omega$, one can understand the maximal amount of breaking expected in the presence of a NP scalar.
@@
Figure 13: Plot of the position of the zero of $P_{2}$ and $P_{4,5}^{\prime}$ versus $C_{9}^{\rm NP}$. The dashed line corresponds to the approximate expression of Eq. 68.

### v2

@@ I Introduction
The purpose of these analyses was to extract Wilson coefficients together with hadronic information directly from the data using an Effective Field Theory (EFT)-based functional form for the transversity amplitudes. These analyses differ in the way that the hadronic contributions are parameterised. On the one hand, Ref. Aaij et al. (2024a) relies on the use of a $z$-expansion to parameterise the non-local charm-loop contributions and includes constraints from $B(B\to K^{*}J/\psi)$ and $B(B\to K^{*}\psi(2s))$ decays. In this analysis, two hypotheses are considered for the data included in the fit: (i) $q^{2}<0$, where constraints coming from the light-cone Operator Product Expansion (OPE) calculation in Ref. Gubernari et al. (2021, 2022) are included; and (ii) $q^{2}>0$, where these constraints are excluded. By contrast, the alternative analysis of Ref. Aaij et al. (2024b) uses a dispersive approach based on the resonance model of Ref. Cornella et al. (2020) and data from the decay $B\to K^{*}\mu^{+}\mu^{-}$ alone. An interesting outcome of both of these analyses is that the data indicate that the charm-loop is insufficient to explain the observed anomalies, and a new physics contribution is required.
@@ II SM predictions for the unbinned observables
While the contributions from local form factors are included in $\mathcal{A}_{V,A}^{\mu}$, non-local quark loop contributions appear in the term $\mathcal{H}^{\mu}$. In the low-$q^{2}$ region, these effects are taken into account by recasting the phenomenological parameterisation provided in Eq. (7.14) of Ref. Khodjamirian et al. (2010). This parameterisation is itself based on the light-cone OPE and LCSR calculations of the non-local form factor in the region $q^{2}\ll 4m_{c}^{2}$, and its corresponding matching to the dispersion relation proposed in the same paper. In practice, this is implemented as a shift in the factorisable charm-quark loop function $Y(q^{2})$ Buras et al. (1994)
@@
where $s_{0}=1$ GeV2. The parameters $a^{\perp,\parallel,0}$, $b^{\perp,\parallel,0}$, and $c^{\perp,\parallel,0}$ are determined by ensuring that they reproduce the phenomenological model of Ref. Khodjamirian et al. (2010) in the $q^{2}$ range from 1 $1\text{ GeV}^{2}$ to 9 $9\text{ GeV}^{2}$. Specific values for these coefficients can be found in Ref. Descotes-Genon et al. (2016). See also Ref. Bordone et al. (2024) for another treatment of long-distance amplitudes.
@@
In the high-$q^{2}$ region several charmonium resonances appear. In order to avoid model-dependence biases we do not attempt to account for non-local contributions in this region. A common approach in the literature is to use quark-hadron duality Shifman (2000); Grinstein and Pirjol (2004); Beylich et al. (2011); Lyon and Zwicky (2014), assuming that when integrated over a large bin the inclusion of these broad charmonium resonances averages to the short-distance contribution. However, this would not be the case in an unbinned analysis. Analyses, such as Ref. Aaij et al. (2024b) use a model of Breit-Wigner amplitudes fitted to the data to try and explicitly include such resonances. Such an approach is inherently model dependent. In our proposed approach, model-independent observables are obtained, such that any estimation of the effect of the wide resonances can be evaluated a posteriori. In particular, should theoretical advances be made concerning the calculation of these resonances, the model-independent data may be reinterpreted.
@@ II.1 Definition of unbinned observables
The uncertainties for the SM predictions are computed using a Monte Carlo sampling approach, which assumes a multivariate Gaussian distribution for the nuisance parameters described earlier, along with variations in the renormalisation scale, masses, and other relevant factors. The structure of the angular observables gives rise to non-linear behaviour, particularly stemming from the normalisation terms. A Gaussian scan of the parameters can therefore lead to non-Gaussian distributions of the observables. This is particularly the case when the normalisation is close to zero.
To extract meaningful uncertainties, we linearise the error propagation. This is effectively achieved by introducing a rescaling factor to the uncertainties associated with all input parameters. These rescaled uncertainties are then used as inputs for the corresponding probability distributions from which we generate the Monte Carlo replicas. Each replica is propagated through the mathematical structure of the observables to compute the full set of observables relevant to the analysis.
The resulting ensemble of observable samples is then used to produce predictions by computing the sample mean and covariance. When evaluating the covariance, the rescaling factor is reintroduced to avoid underestimating the overall uncertainty. We observe that a rescaling factor of $1/2$ is already sufficient to achieve this behaviour, and increasing it further does not lead to significant changes in the resulting distribution of observables. Nevertheless, our default choice for the rescaling factor is $1/3$.
For further details on this procedure, we refer the reader to Section 4.1 of Ref. Descotes-Genon et al. (2016). The use of the full distribution of observables for the characterisation of relevant statistical metrics will be discussed in a forthcoming publication by some of the authors Capdevila and Scaffidi ((in preparation).
@@
The SM prediction for the optimised observables $P_{i}$ and $F_{L}$ are shown graphically in Figs. 1 and 2. The $P_{i}^{\rm CP}$ are, to a high accuracy, very close to zero and we do not include them here. For completeness we also provide the non-optimised observables $S_{i}$ and the branching fraction in Appendix A. The unbinned definitions of the $S_{i}$ observables can be easily derived following the same procedure as in Eq. (12). Some remarks are in order here:
@@
The next issue to address is how to approach unbinned data. Here there are two main options: i) either one parameterises the amplitudes with some well-motivated, but model-dependent, theoretical guidance and computes the observables from these amplitudes or ii) one determines the observables directly. In the former approach, LHCb data are fitted using a model for the amplitudes that incorporates both local and non-local contributions. From this fit, parameters associated with the non-local terms, as well as the Wilson coefficients that characterise the local contributions, are extracted. The extraction of these parameters depends on the underlying assumptions, such as the inclusion of potential contributions from NP in the form of scalars, tensors, or additional operators beyond the SM, the modelling of hadronic effects, and the treatment of the variables $m_{K\pi}$ and $q^{2}$. Once the amplitudes are determined, the observables can a posteriori be reconstructed. There are different analyses in the literature parameterising the amplitudes Aaij et al. (2024a, b). In contrast, the latter approach, extracting observables directly from the distributions, avoids this model dependence. This method allows us to obtain information without imposing a priori assumptions, enabling reinterpretation with new theoretical understanding and various tests of the internal consistency of the experimental analysis through the use of the observables themselves. In the following subsections II.2 and II.3 we define the tests that a consistent theoretical model of the amplitudes should fulfill. In section III, we describe a model-independent method to determine unbinned observables using a moments approach.
@@ II.2 Testing a theoretical model of the amplitudes: set of conditions
If a parameterisation of the amplitudes does not respect the conditions, for instance, by fixing some degrees of freedom with theory constraints in a way that is inconsistent with the data, this can manifest itself as the inability to reproduce one or more observables. In order to remedy this, more degrees of freedom must be included or some of the fixed degrees of freedom must be allowed to vary.
@@
In the presence of scalar NP, the last two conditions in Eq. (LABEL:nJ) are affected and also the last condition in Eq. (17). In both cases, the left hand side (LHS) of these conditions has to be modified. In a similar way, one can write a corresponding set of equations for the CP conjugate $J$ and $n_{i}$.
@@ II.3 Interpretation of conditions
In order to test a given amplitude parameterisation using the conditions in Eq. (II.2) and Eq. (17), on the LHS of these expressions one should insert the parameterisation of the amplitudes as a function of the relevant free parameters (phases and moduli); and on the RHS, experimental values for the coefficients $J$. A fit for the free parameters can then be used to check if it is possible to reproduce the RHS of each equation.
@@
    A fit to data using only $q^{2}>0$ information but adding lattice calculations. The fit for the observables in Ref. Aaij et al. (2024a) reproduces the binned data very well. We therefore take this fit for the observables as the “data”, i.e., the RHS of the conditions in Eq. (II.2) and Eq. (17). The results of this fit are depicted in blue in Fig. 3 and in Fig. 15 in Appendix. B.
@@
    A fit to data using, in addition, the $q^{2}<0$ information. This fit is highly impacted by theory points at $q^{2}<0$ and we reinterpret it as the LHS of the conditions. The results of this fit are depicted in red in Fig. 3 and in Fig. 15.
@@
The tension observed in condition 4 is linked to the tension observed in the comparison between ${\rm Im}A_{\|}$ for $q^{2}>0$ and $q^{2}<0$ in Ref. Aaij et al. (2024a). To a much lesser extent there is also a tension in ${\rm Im}A_{\perp}$, as discussed below.
@@
${A}_{\lambda}^{L,R}$ $=$ ${\cal N}_{\lambda}\left(\left[C_{9}\mp C_{10}\right]{\cal F}_{\lambda}(q^{2})+\frac{2m_{b}M_{B}}{q^{2}}[C_{7}{\cal F}_{\lambda}^{T}(q^{2})\right.$ (22)
$\left.-16\pi^{2}\frac{M_{B}}{m_{b}}{\cal H}_{\lambda}]\right)$
@@ III A new method to determine unbinned observables: A “continuous” moments method
Here we introduce a novel method to extract the full set of model-independent angular observables continuously in $q^{2}$ and $m_{K\pi}$. We do this by extending the previous binned method-of-moments technique Beaujean et al. (2015a) into an unbinned one. This is achieved with the use of a $k$-nearest neighbours (knn) algorithm Cover and Hart (1967). The knn with uniform weighting is exactly the definition of the normalised moment (the moment divided by the total weight), $M_{i}$, but instead of being binned, it is really a reflection of the density of the data in the locale of a given phase-space point. The knn takes the k nearest events in a region of $q^{2}$, $m_{K\pi}$ or both for the given point at which the data are being evaluated. Each of the $k$ events has a value of $\cos\theta_{K}$, $\cos\theta_{\ell}$ and $\phi$, and so for each the moment is calculated, i.e. each event has a value of $m_{5}=\sin 2\theta_{K}\sin\theta_{\ell}\cos\phi$. The knn algorithm sums the relevant moment ($m_{i}$) and divides by $k$ to calculate the average moment in that region. Thus one can pick an arbitrary point in $q^{2}$, $m_{K\pi}$, or $(q^{2},m_{K\pi})$, and estimate the value of the angular moment there.
@@
With this unbinned method, different points in the phase space are statistically correlated with each other. The degree of correlation depends on the number of neighbours in the knn, the smoothing and how far apart the points are. When using the results of this method in further analysis for the underlying physics parameters, knowing these correlations is vital. Consequently, if the correlations are well understood, the precise choice of the number of neighbours or smoothing strength is to some extent a matter of aesthetics. It is recommended that experimental analyses publish unsmoothed distributions, such that theorists are free to make their own choices in their subsequent analyses. Few neighbours leads to a large apparent uncertainty at a single point but nearby points are relatively independent and so give equal statistical power. Conversely, a large number of neighbours reduces the apparent uncertainty but leads to large statistical correlations with neighbouring regions. What really matters is choosing the analysis parameters that maximise the sensitivity to what is being measured, and in particular allows for the resolution of the phase-space dependence of the observables. Care must be taken, as using too many neighbours can lead to a poor description of the true variables in regions where they vary quickly. However, for some analyses, such as finding the zero-crossing points in $q^{2}$ (discussed in the following sections), one may choose a larger number of neighbours, as the zeros lie in a $q^{2}$ region where the observables of interest are expected to vary smoothly.
The output of the method is a large grid of the values of the angular observables at each of the sampled points in the variables of interest ($q^{2}$, $m_{K\pi}$ or both), with the size of the grid depending on the number of sampled points, $N$. The choice of $N$ should be large enough such that all of the experimental data are exploited (i.e. there are no regions in the variable of interest that are not covered by the $knn$ output) but not so large that the output becomes unwieldy. By evaluating the correlations between observables and points in e.g. $q^{2}$, one has a $MN\times MN$ covariance matrix, where $M$ is the number of angular observables (11 in the P-wave with massive leptons). This matrix is positive definite and sparse as regions of the variables that are separated by any distance will have negligible correlations. Inverting this matrix for subsequent fits of the physics parameters is necessarily computationally expensive but need only be done once. The sparse nature of the matrix implies that numerical methods may be used to approximate the inverse, such as conjugate gradient descent Hestenes and Stiefel (1952).
@@
Previous experimental analyses have assumed that the leptons are massless, reducing the number of angular terms to be measured and so simplifying the task. However, it is found that the effects of lepton mass in the muon mode are non-negligible even at relatively large values of $q^{2}$ ($\sim 4\,\,\text{GeV}^{2}$), as shown in Fig. 5 (see also Appendix C for a definition of a simpler lepton mass modified distribution that avoids the conflicting terms $J_{1}^{s}$ and $J_{1}^{c}$). In the absence of S-wave contributions, the full set of angular observables may be extracted with the unbinned moments, fully accounting for the effects of the lepton masses. The observables $S_{1}^{s}$, $S_{1}^{c}$, $S_{2}^{s}$, $S_{2}^{c}$ do not have orthogonal angular functions and must be extracted from a linear combination of the corresponding moments
@@ IV.1.1 Geometrical tests associated with the structure of the observables
It is important to notice when comparing to data that the LHCb collaboration uses a different convention for the observables than the original definition in Ref. Descotes-Genon et al. (2013b). These differing conventions can be interchanged using,
@@
The terms $\omega_{r,i}$ contain the scalar contributions. If the signs of $P_{5}^{\prime}$ and $\omega_{r}$ are the same the bound in Eq. (39) may be violated in an analysis measuring observables, indicating the presence of a scalar NP contribution according to Eq. (47). Similarly, if the signs of $P_{6}^{\prime}$ and $\omega_{i}$ are the same, the bound in Eq. (35) may also be violated according to Eq. (48). All other bounds are unaffected by the presence of a scalar.
@@ IV.1.2 Tests based on the non-trivial relation between the observables
At this point there are two options: 1) to test directly the relation in Eq. (52), which is given in terms of $J_{i}$; or 2) to re-express it in terms of the $P_{i}$ and $P_{i}^{\rm CP}$ observables, as shown below. The latter is implemented by inverting the definitions in Eq. (12). There is a minor difference between the two options, because in the first case a breaking of Eq. (52) can only be due to an internal problem of the experimental analysis, or to a large new scalar contribution. In the second case, if we use the expression in terms of $P_{i}$ and $P_{i}^{\rm CP}$ and then assume that all $P_{i}^{\rm CP}$ are negligible, we introduce a third possible source of breaking if there exists large NP weak phases in Nature. In any case, this relation is always a very efficient cross-check of an experimental analysis.
We have explicitly checked that, up to machine precision, the relation of Eq. (52) using the $J$’s obtained from the fit of Ref. Aaij et al. (2024a) is fulfilled, considering either the $q^{2}<0$ or $q^{2}>0$ cases. Since the $J_{i}$ are obtained from an amplitude analysis where no scalar contribution is introduced, the relation in Eq. (52) must, by construction, be fulfilled. However, this implies nothing about the presence or absence of a scalar amplitude.
In order to express Eq. (52) in terms of optimised observables we use the definitions of these observables in Eq. (12). One can then write down one relation for $\bar{P_{i}}=P_{i}+P_{i}^{\rm CP}$ and one for $\hat{P_{i}}=P_{i}-P_{i}^{\rm CP}$. Taking the example of $P_{2}$, we can write an equation for $\bar{P}_{2}$ and one for $\hat{P}_{2}$ as a function of all the other observables. The corresponding equations are Matias and Serra (2014)
$\bar{P}_{2}=P_{2}+P_{2}^{\rm CP}=+\frac{1}{2{\bar{k}_{1}}}\bigg{[}({\bar{P}_{4}^{\prime}}{\bar{P}_{5}^{\prime}}+\delta_{1})+\frac{1}{\beta}\sqrt{(-1+{\bar{P}_{1}}+{\bar{P}_{4}^{\prime 2}})(-1-{\bar{P}_{1}}+\beta^{2}{\bar{P}_{5}^{\prime 2}})+\delta_{2}+\delta_{3}{\bar{P}_{1}}+\delta_{4}{\bar{P}_{1}}^{2}}\bigg{]},$ (53)
@@
$\hat{P}_{2}=P_{2}-P_{2}^{\rm CP}=+\frac{1}{2{\hat{k}_{1}}}\bigg{[}({\hat{P}_{4}^{\prime}}{\hat{P}_{5}^{\prime}}+\hat{\delta}_{1})+\frac{1}{\beta}\sqrt{(-1+{\hat{P}_{1}}+{\hat{P}_{4}^{\prime 2}})(-1-{\hat{P}_{1}}+\beta^{2}{\hat{P}_{5}^{\prime 2}})+\hat{\delta}_{2}+\hat{\delta}_{3}{\bar{P}_{1}}+\hat{\delta}_{4}{\hat{P}_{1}}^{2}}\bigg{]},$ (54)
@@
where $\Omega=\frac{1}{\beta}\sqrt{(-1+P_{1}+{P_{4}^{\prime}}^{2})(-1-P_{1}+\beta^{2}{P_{5}^{\prime}}^{2}})$. For illustration, we show in Fig. 11 the expression in Eq. (55) rewritten in quadratic form using the results from one pseudo-experiment.
As discussed at the beginning of this subsection IV.1, in the presence of a NP scalar the non-trivial relation between the observables is broken. Here we estimate, for the first time, the amount of breaking of the relation due to the presence of a scalar by combining theory inputs with data on the branching fraction of ${B_{s}\to\mu^{+}\mu^{-}}$. If we allow for the presence of a scalar, Eq. (53) and Eq. (54) can be generalised by introducing the following substitutions in Eq. (53)
@@
are the CP counterparts of the $\omega_{r,i}$ defined in Eq. (49) and Eq. (50), with $N$ given by Eq. (51). By expanding the terms in $\omega$, one can understand the maximal amount of breaking expected in the presence of a NP scalar.
@@
Figure 13: Plot of the position of the zero of $P_{2}$ and $P_{4,5}^{\prime}$ versus $C_{9}^{\rm NP}$. The dashed line corresponds to the approximate expression of Eq. (68).